When a person leaves his home for sightseeing by his car, the meter reads 12352 km. When he returns home after two hours the reading is 12416 km. (a) What is the average speed of the car during this period? (b) What is the average velocity?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. (a) Total distance covered 12416 – 12352 = 64 km in 2 hours. Speed = 64/2 = 32 km/h (b) As he returns to his house, the displacement is zero. Velocity. = (displacement / time) = 0 (zero).